A plane flies 720 mi against a steady 30-mi/h headwind and then returns to the same point with the wind. If the entire trip takes 10h, what is the plane's speed in still air?
Answer by checkley71(8403)

10=720/(X-30)+720/(X+30)
10=[720(X+30)+720(X-30)/(X-30)(X+30)]
10=[720X+2160+720X-2160]/(X^2-900) NOW CROSS MULTIPLY
10X^2-9000=1440X
10X^2-1440X-9000=0
10(X^2-144-900)=0
10(X-150)(X+6)=0
X-150=0
X=150 FOR THE SPEED OF THE AIRPLANE IN STILL AIR.
PROOF
10=(720/120+720/180)
10=6+4
10=10